OpenStudy (anonymous):

Use elimination to find the solution to the system of equations. 9x+2y=-45 2x+6y=-10

OpenStudy (anonymous):

@iGreen

OpenStudy (irishboy123):

**elimination** \(9x+2y=-45 💙 \\ 2x+6y=-10 💚\) \(27x+6y=-135 💙 [i.e. \times 3]\\ 2x+6y=-10 💚 \) \(27x+6y=-135 💙\\ -2x-6y=+10 💚 [i.e. \times -1]\) now what is 💙 + 💚 ????

OpenStudy (anonymous):

145?

OpenStudy (anonymous):

@Irishboy123

OpenStudy (irishboy123):

\(\{27x+6y=-135\} \\ + \\\{ -2x-6y=+10\}\)